To raise the temperature of 75.0 g of a particular metal by 1.50 celcius reuires 107 cal of heat. What approximate atomic mass of tge metal? what is the metal?
1
Expert's answer
2019-09-04T03:23:27-0400

"q=mCdt"

where "m,c" and "dt" represent mass of metal,specific heat capacity and change in temperature respectively.

By substituting the values of the variables above,

"107=75 \times C \times 1.5 \space Cal"

"C=\frac{107 }{75 \times 1.5} \space Cal/gm-\degree C" "=0.951 \space Cal/gm-\degree C" "=0.951 \times 4.184 \approx 4 J/gm-\degree C"

The highest specific heat capacity of a metal is 3.56 (in same units) which is for lithium.

So,the metal as illustrated in the question is not practically possible.

This can be water"(C=4.2)" or some other liquid with specific heat capacity 4.